# First-place Solution for Streetscape Shop Sign Recognition Competition

### v1

@@ 1.1 Background
Storefront signboards  [72] are crucial for extracting essential information [55, 89, 77, 88, 90, 91] from street view imagery, making them significant research subjects in the realm of computer vision. The technology used for character recognition on these signboards integrates visual data processing with practical applications  [24, 25, 60, 62, 63], such as map navigation [46], smart city planning  [65, 64, 66], and commercial value analysis, presenting substantial research and commercial opportunities  [27, 9, 8, 31]. Participants in this competition are charged with the development of an advanced shop signboard recognition and analysis system [1, 57, 86, 58, 84, 78, 79, 82, 80, 81, 83, 87]. Using street view photos and labels provided by the organizers, We face the challenge of precisely detecting shop signboards, analyzing their layouts, and recognizing shop names within the context of natural street view imagery. The competition is centered around natural street scenes [4, 3, 51, 19, 56, 49, 26, 52], specifically targeting the storefront signboards and their textual content, with a focus on daytime images only. The given input is a street view image, and the required outputs are the locations of the shop signs (limited to visible signboards) and the names of the shops as depicted on these signboards. The primary objective of this competition is to significantly improve the accuracy of signboard detection [48, 50, 47, 76, 92, 23] and the precision of the shop names recognized [6, 7, 94, 22, 85, 36, 34, 33, 38, 37, 40, 44, 43, 45] in the test set.
@@ 1.3.1 Dataset
Three distinct datasets have been provided for this challenge: ”Store sign and store name data,” ”Signboard detection data,” and ”Signboard OCR data.” The evaluation set will adhere to the format of the ”Store sign and store name data.” Below is an overview of each dataset: Store sign and store name data. This dataset features street view images with potential multiple storefronts per image. Each storefront may display several signboards, some of which are unrelated to stores, and some that are explicitly store signboards. There is a possibility that some images do not contain any store signboards. Each image in the dataset is labeled with the valid quadrilateral box of the store signboard and its corresponding store name. Valid signboards are defined as having a complete and clear shape, minimal occlusion [35, 39, 32, 42, 41], and displaying a specific store name. The training set comprises approximately 5,000 images, while the test set contains about 500 images. In the test set, store names are required to have at least two Chinese characters and must not include English characters or special symbols [13, 12, 11]. The label file is in text format, where each line specifies the coordinates of the four vertices of the signboard, arranged clockwise, and the corresponding store name. Sample images from the dataset are depicted in Fig. 1.
@@ 2.2 Signboard Text Detection & Key Information Extraction (KIE)
In the stage of text detection [71, 30, 75, 96] and signboard text extraction phase, the two tasks are combined for end-to-end training due to the commonality of text features. The pipeline of the model is shown in Fig. 5. First, the model uses a two-stage detector to detect the location of the text and extract both the positional embedding and the image embedding of the text. A graph neural network is then used to integrate multimodal features and distinguish which belong to store signboard texts. The end-to-end training paradigm, which is inspired by the former work [71], allows for more powerful and effective features to be learned, thereby improving the accuracy of both tasks simultaneously. In addition, we use a reinforcement learning-based method called the BoxDQN [69], which adjusts the shape of the text box to find the optimal box for the recognition model, thus improving the accuracy of end-to-end recognition.
@@ 2.3 Text Recognition
In the text recognition phase, we judge the direction of the text based on the aspect ratio of the text box, and then train two text recognition models [67, 15, 95, 54, 70] for horizontal and vertical text, respectively. The baseline of the text recognition model we use is SAR [li2019show]. Based on it, we make some improvements. (1) upgrading the backbone to VIT [14], (2) changing the unidirectional LSTM to bidirectional LSTM [20], integrating LSTM and 2D attentional features in parallel, and providing richer contextual information. (3) using multiple self-supervised training, and (4) adding Center Loss to boost the discrimination of morphological close word recognition features. Two self-supervised pre-training methods that contribute significantly are described below. The first one is a pre-training method based on Sequential Contrast Learning [2, 68], as shown in Fig. 6. In detail, we perform some augmentation on the text image to get a positive sample pair and then select another text image as a negative sample pair. The model extracts the sequence features using an encoder and then learns from the positive and negative sample pairs based on the sequence. The second one is masking pre-training based on stroke information and semantic context. As shown in Fig. 7, it is an improvement based on MAE [18], where different masks are set for stroke features and semantic features to predict and learn the masks. In this way, the mural information and textual information of the text can be modeled simultaneously to improve the feature representation of the model. Using these two different self-supervised pre-training methods, we can obtain an extractor with strong modeling ability, thus improving the performance of text recognition.
@@ 2.5 Results
Overall, our approaches can be found in Tab. 1. In this competition, we incorporate three techniques that have become popular in recent years: multimodal feature fusion, large-scale self-supervised training, and Transformer-based large model. In addition, depending on the properties of the data provided, we add some illuminating methods such as BoxDQN based on reinforcement learning and text rectification, which achieve promising performance.
@@ 3.1 Evaluation Metric
In order to support domestic machine learning development, the competition uses the Jittor[21] framework to provide the baseline code of the competition. Jittor is a high-performance deep learning framework based on instant compilation and meta operators. The entire framework integrates powerful Op compilers and tuners with instant compilation to generate customized high-performance code for your model. Jittor also contains a rich high-performance model library, covering image recognition, detection, segmentation, generation, differentiable rendering, geometry learning, reinforcement learning, etc.
The evaluation metrics of the competition are the accuracy and recall rates of store signboard identification. Only when the location of the detection frame of the shop signboard is accurate and the identification of the shop name line in the signboard is correct, the identification is defined to be accurate. The formula for calculating the recall rate of store sign recognition is the ratio of the number of correctly recognized store signs to the true number of store signs. The calculation formula of the accuracy rate of store sign recognition is the ratio of the number of correctly recognized store signs to the total number of recognized store signs. The final score is the F-score of accuracy rate and recall rate, calculated as $2*accuracy*recall/(accuracy+recall)$.
@@ 3.2 Competition Results
There are many challenges in street view store signboard recognition tasks, such as various modes and contents of store signboards, geometric distortion and truncation in image shooting, and limited training samples. Our team, with its profound accumulation in the OCR field for many years, has better analyzed the difficulties and proposed a reasonable algorithm framework according to the characteristics of the competition, and won first place in the competition. The overall framework divides the competition task into four stages: signboard detection, signboard text detection, text recognition, and reading sequence prediction. The main highlights are” signboard text detection and key information extraction”, ”multi-modal modeling &GNN ”,” text self-supervised pre-training” and ”VIT model”. This method effectively integrates multi-dimensional information such as image features and text position distribution for multi-modal recognition and reasoning. In the case of limited samples, self-supervised pre-training, and data augmentation methods are used to significantly improve the performance in a short period of time.
@@ 4 Conclusion
The profound development of deep learning has catalyzed significant advancements in the field of computer vision. Previously confined to elementary tasks such as detection, segmentation, and classification, the understanding of images has now expanded to encompass scene understanding, multi-modal recognition, relational reasoning, and comprehensive end-to-end image interpretation. In the specific context of text recognition in street view store signboards, computer vision research has identified several key evolving trends: 1) Adoption of Semi-supervised or Unsuspected Learning: Large-scale pre-training techniques that rely on semi-supervised or unsupervised learning have become prevalent. These methods focus on deciphering the subtle relationships and intrinsic connections within roadway data, providing robust pre-training representations that bolster downstream perception tasks. This approach not only reduces the necessity for finely labeled data but also enhances the model’s generalization capabilities and diminishes the costs associated with research and development. 2) Multi-dimensional, Multi-modal Feature Analysis: The use of sophisticated multi-modal, multi-task object relationship analysis based on multi-dimensional features has proven effective in elevating recognition accuracy. 3) Enhanced Scene and Quality Recognition: Advanced scene recognition and quality assessments are utilized to sift through and exclude non-relevant data, ensuring that only pertinent information is processed. 4) Progress in End-to-End Image Understanding: Rapid advancements in end-to-end methodologies are setting the stage for the emergence of real-world map auto-generation technologies. These sophisticated systems integrate targeted content analysis, detailed scene evaluation, and precise location restoration. Looking ahead, these innovations are poised to significantly impact various sectors, playing a pivotal role in navigation systems, intelligent recommendations, smart city development, and autonomous driving technologies. In the near future, with the development of multimodal large models  [5, 73, 93, 16, 74], end-to-end signature text reading becomes possible with a higher ceiling.

### v2

@@ 1.1 Background
Storefront signboards  [74] are crucial for extracting essential information [56, 92, 79, 91, 93, 94] from street view imagery, making them significant research subjects in the realm of computer vision. The technology used for character recognition on these signboards integrates visual data processing with practical applications  [25, 26, 61, 63, 64], such as map navigation [47], smart city planning  [66, 65, 67], and commercial value analysis, presenting substantial research and commercial opportunities  [28, 9, 8, 32]. Participants in this competition are charged with the development of an advanced shop signboard recognition and analysis system [1, 58, 89, 59, 86, 80, 81, 84, 82, 83, 85, 90]. Using street view photos and labels provided by the organizers, we face the challenge of precisely detecting shop signboards, analyzing their layouts, and recognizing shop names within the context of natural street view imagery. The competition is centered around natural street scenes [4, 3, 52, 19, 57, 50, 27, 53], specifically targeting the storefront signboards and their textual content [70, 76], with a focus on daytime images only. The given input is a street view image, and the required outputs are the locations of the shop signs (limited to visible signboards) and the names of the shops as depicted on these signboards. The primary objective of this competition is to significantly improve the accuracy of signboard detection [49, 51, 48, 78, 95, 24] and the precision of the shop names recognized [6, 7, 97, 23, 87, 37, 35, 34, 39, 38, 41, 45, 44, 46] in the test set.
@@ 1.3.1 Dataset
Three distinct datasets have been provided for this challenge: ”Store sign and store name data,” ”Signboard detection data,” and ”Signboard OCR data.” The evaluation set will adhere to the format of the ”Store sign and store name data.” Below is an overview of each dataset: Store sign and store name data. This dataset features street view images with potential multiple storefronts per image. Each storefront may display several signboards, some of which are unrelated to stores, and some that are explicitly store signboards. There is a possibility that some images do not contain any store signboards. Each image in the dataset is labeled with the valid quadrilateral box of the store signboard and its corresponding store name. Valid signboards are defined as having a complete and clear shape, minimal occlusion [36, 40, 33, 43, 42], and displaying a specific store name. The training set comprises approximately 5,000 images, while the test set contains about 500 images. In the test set, store names are required to have at least two Chinese characters and must not include English characters or special symbols [13, 12, 11, 88, 20]. The label file is in text format, where each line specifies the coordinates of the four vertices of the signboard, arranged clockwise, and the corresponding store name. Sample images from the dataset are depicted in Fig. 1.
@@ 2.2 Signboard Text Detection & Key Information Extraction (KIE)
In the stage of text detection [73, 31, 77, 99] and signboard text extraction phase, the two tasks are combined for end-to-end training due to the commonality of text features. The pipeline of the model is shown in Fig. 5. First, the model uses a two-stage detector to detect the location of the text and extract both the positional embedding and the image embedding of the text. A graph neural network is then used to integrate multimodal features and distinguish which belong to the store signboard texts. The end-to-end training paradigm, which is inspired by the former work [73], allows for more powerful and effective features to be learned, thereby improving the accuracy of both tasks simultaneously. In addition, we use a reinforcement learning-based method called the BoxDQN [71], which adjusts the shape of the text box to find the optimal box for the recognition model, thus improving the accuracy of end-to-end recognition.
@@ 2.3 Text Recognition
In the text recognition phase, we judge the direction of the text based on the aspect ratio of the text box, and then train two text recognition models [68, 15, 98, 55, 72] for horizontal and vertical text, respectively. The baseline of the text recognition model we use is SAR [li2019show]. Based on it, we make some improvements. (1) upgrading the backbone to VIT [14], (2) changing the unidirectional LSTM to bidirectional LSTM [21], integrating LSTM and 2D attentional features in parallel, and providing richer contextual information. (3) using multiple self-supervised training, and (4) adding Center Loss to boost the discrimination of morphologically close word recognition features. Two self-supervised pre-training methods that contribute significantly are described below. The first one is a pre-training method based on Sequential Contrast Learning [2, 69], as shown in Fig. 6. In detail, we perform some augmentation on the text image to get a positive sample pair, and then select another text image as a negative sample pair. The model extracts the sequence features using an encoder and then learns from the positive and negative sample pairs based on the sequence. The second one is masking pre-training based on stroke information and semantic context. As shown in Fig. 7, it is an improvement based on MAE [18], where different masks are set for stroke features and semantic features to predict and learn the masks. In this way, the mural information and textual information of the text can be modeled simultaneously to improve the feature representation of the model. Using these two different self-supervised pre-training methods, we can obtain an extractor with strong modeling ability, thus improving the performance of text recognition.
@@ 2.5 Results
Overall, our approaches can be found in Tab. 1. In this competition, we incorporate three techniques that have become popular in recent years: multimodal feature fusion, large-scale self-supervised training, and a Transformer-based large model. In addition, depending on the properties of the data provided, we add some illuminating methods such as BoxDQN based on reinforcement learning and text rectification, which achieve promising performance.
@@ 3.1 Evaluation Metric
In order to support domestic machine learning development, the competition uses the Jittor[22] framework to provide the baseline code of the competition. Jittor is a high-performance deep learning framework based on instant compilation and meta operators. The entire framework integrates powerful Op compilers and tuners with instant compilation to generate customized high-performance code for your model. Jittor also contains a rich, high-performance model library, covering image recognition, detection, segmentation, generation, differentiable rendering, geometry learning, reinforcement learning, etc.
The evaluation metrics of the competition are the accuracy and recall rates of store signboard identification. Only when the location of the detection frame of the shop signboard is accurate and the identification of the shop name line in the signboard is correct, is the identification defined to be accurate. The formula for calculating the recall rate of store sign recognition is the ratio of the number of correctly recognized store signs to the true number of store signs. The calculation formula of the accuracy rate of store sign recognition is the ratio of the number of correctly recognized store signs to the total number of recognized store signs. The final score is the F-score of accuracy rate and recall rate, calculated as $2*accuracy*recall/(accuracy+recall)$.
@@ 3.2 Competition Results
There are many challenges in street view store signboard recognition tasks, such as various modes and contents of store signboards, geometric distortion and truncation in image shooting, and limited training samples. Our team, with its profound accumulation in the OCR field for many years, has better analyzed the difficulties and proposed a reasonable algorithm framework according to the characteristics of the competition, and won first place in the competition. The overall framework divides the competition task into four stages: signboard detection, signboard text detection, text recognition, and reading sequence prediction. The main highlights are” signboard text detection and key information extraction”, ”multi-modal modeling &GNN ”,” text self-supervised pre-training”, and ”VIT model”. This method effectively integrates multi-dimensional information such as image features and text position distribution for multi-modal recognition and reasoning. In the case of limited samples, self-supervised pre-training and data augmentation methods are used to improve the performance significantly in a short period of time.
@@ 4 Conclusion
The profound development of deep learning has catalyzed significant advancements in the field of computer vision. Previously confined to elementary tasks such as detection, segmentation, and classification, the understanding of images has now expanded to encompass scene understanding, multi-modal recognition, relational reasoning, and comprehensive end-to-end image interpretation. In the specific context of text recognition in street view store signboards, computer vision research has identified several key evolving trends: 1) Adoption of Semi-supervised or Unsupervised Learning: Large-scale pre-training techniques that rely on semi-supervised or unsupervised learning have become prevalent. These methods focus on deciphering the subtle relationships and intrinsic connections within roadway data, providing robust pre-training representations that bolster downstream perception tasks. This approach not only reduces the necessity for finely labeled data but also enhances the model’s generalization capabilities and diminishes the costs associated with research and development. 2) Multi-dimensional, Multi-modal Feature Analysis: The use of sophisticated multi-modal, multi-task object relationship analysis based on multi-dimensional features has proven effective in elevating recognition accuracy. 3) Enhanced Scene and Quality Recognition: Advanced scene recognition and quality assessments are utilized to sift through and exclude non-relevant data, ensuring that only pertinent information is processed. 4) Progress in End-to-End Image Understanding: Rapid advancements in end-to-end methodologies are setting the stage for the emergence of real-world map auto-generation technologies. These sophisticated systems integrate targeted content analysis, detailed scene evaluation, and precise location restoration. Looking ahead, these innovations are poised to significantly impact various sectors, playing a pivotal role in navigation systems, intelligent recommendations, smart city development, and autonomous driving technologies. In the near future, with the development of multimodal large models  [5, 75, 96, 16, 76], end-to-end signature text reading becomes possible with a higher ceiling.